# SEAL: Towards Safe Autonomous Driving via Skill-Enabled Adversary Learning for Closed-Loop Scenario Generation

### v2

@@ Abstract
Verification and validation of autonomous driving (AD) systems and components is of increasing importance, as such technology increases in real-world prevalence. Safety-critical scenario generation is a key approach to robustify AD policies through closed-loop training. However, existing approaches for scenario generation rely on simplistic objectives, resulting in overly-aggressive or non-reactive adversarial behaviors. To generate diverse adversarial yet realistic scenarios, we propose SEAL, a scenario perturbation approach which leverages learned objective functions and adversarial, human-like skills. SEAL-perturbed scenarios are more realistic than SOTA baselines, leading to improved ego task success across real-world, in-distribution, and out-of-distribution scenarios, of more than 20%. To facilitate future research, we release our code and tools: https://navars.xyz/seal/
@@ I Introduction
Therefore, in this paper, we propose and evaluate a method for Skill-Enabled Adversary Learning (SEAL), which yields significantly improved downstream ego behavior, in closed-loop training with safety-critical scenario generation. Our method addresses the identified limitations in prior art by introducing two novel components, as shown in LABEL:fig:overview. First, we introduce a learned objective function to anticipate how a reactive ego agent will respond to a candidate adversarial agent behavior. We quantify both collision closeness and induced ego behavior deviation, thus providing a broadened understanding of safety criticality. Second, we develop a skill-enabled, reactive adversary policy; in particular, inspired by human cognition, we leverage a hierarchical framework that is akin to how humans operate vehicles [27] and we create an adversarial prior that selects human-like skill primitives to increase criticality while maintaining realism.
Furthermore, we argue that safety-critical scenario generation should be evaluated based on behavior realism and usefulness for ego policy improvement, not just induced criticality. Prior work often assesses ego policies on generated scenarios where safety-critical behavior remains effectively in-distribution with respect to training data and heuristic perturbations [1, 28]. To address this, we build on recent scenario characterization work, SafeShift [18], to identify real (non-generated) but safety-relevant scenarios, enabling a more realistic, out-of-distribution evaluation. While in-distribution performance is informative, real-world performance on challenging scenes is ultimately most important.
@@
    We propose two novel techniques for safety-critical perturbation: (i) a learned objective function, to select candidate trajectories; and (ii) an adversarial skill-based, reactive policy for more realism in adversary behavior.
@@
    We design an improved evaluation setting for closed-loop training, utilizing real-world safety-relevant scenarios in contrast to just in-distribution generated scenes.
@@ II-A Scenario Generation in Autonomous Driving
Approaches for generating scenarios that reproduce the distribution of normal driving behavior have been extensively explored. Some methods ensure the diversity of generated traffic behavior [29, 30], while others aim for controllability through rule-based or language-driven specifications [31, 32, 33]. However, due to the rarity of safety-critical events in recorded data [17, 18, 19], other approaches have focused on directly generating corner-case scenarios by injecting adversarial behaviors. Earlier works in safety-critical scenario generation relied on gradient-based optimization approaches with access to vehicle dynamics [28, 25, 34], a limitation in model-free settings. Other methods, such as diffusion-based approaches [26, 35], are compute-intensive and impractical to be used in a closed-loop manner. Efficient methods like CAT [1] and GOOSE [2], which leverage trajectory prediction priors and reinforcement learning (RL) respectively, prioritize simple collision objectives and are non-reactive to the ego agent. In contrast, our approach efficiently generates reactive, nuanced adversarial behavior, providing a stronger training signal and improving closed-loop agent performance.
@@ II-B Robust Training and Evaluation in Autonomous Driving
Out-of-distribution evaluation has been well-explored in AD trajectory prediction [18, 42, 43, 16], but these approaches often aim to characterize an entire scene without focusing on a single ego driver or identifying a specific adversary. In AD control tasks, some prior work has explored out-of-distribution settings, such as CARNOVEL [44, 45], which tests unseen scenario types like roundabouts. Additionally, Lu et al. [46] evaluate across real-world scenarios of various difficulty levels, but do not hold out the hardest scenes during training. Our approach thus addresses this gap by offering a more comprehensive and rigorous evaluation, across a wide set of adversarial and real-world scenarios.
@@ III Preliminaries
Scenario Perturbation Task: For this task, $K$ re-simulations of a base scenario $\mathbf{S}$ are performed as episodes, where agents start from the same state as the base scenario and follow a behavior prescribed by some policy (i.e., a reactive policy or predefined trajectory), which may be different than their original trajectory. Let $\widetilde{\mathbf{X}}^{(k)}$ represent the observed trajectories in the $k$-th re-simulation of $\mathbf{S}$. The perturbation-based safety critical scenario generation task is thus assigning behaviors to roll-out for all non-ego agents, conditioned on the base scenario $\mathbf{S}$ and $K$ previous episodes, $\{\widetilde{\mathbf{X}}^{(k)}\ |\ k\in\{1,2,...,K\}\}$, such that the resulting $\widetilde{\mathbf{X}}^{(K+1)}$ satisfies some specified desired properties of criticality. Importantly, we treat the ego agent’s behavior as a black box: while we are able to observe previous behavior as $\widetilde{\mathbf{X}}^{K}_{\text{ego}}$, we have no access to the model or any privileged information on ego’s decision-making process.
@@ IV Approach: Skill-Enabled Adversary Learning for Scenario Generation
To increase the criticality while maintaining the realism of a scenario, we propose the Skill-Enabled Adversary Learning (SEAL) approach for perturbation-based scenario generation. We share a similar insight to CAT [1] of utilizing a probabilistic trajectory prediction model, $\pi_{\text{gen}}$, to produce diverse candidate future paths that an adversary could take, conditioned on $\mathbf{S}$. However, there are several challenges with using this previously-explored approach of directly providing a future path for the adversary to follow from $\pi_{\text{gen}}(\mathbf{S})$:
  1. 1.
    Focusing solely on inducing collisions neglects a broader understanding of safety criticality, such as forcing hard braking maneuvers, swerving, or other such deviations from normal ego-driving behavior.
  2. 2.
    Having an adversary follow a predefined path prevents reactivity to an ego agent’s decision making.
  3. 3.
    The heuristically-selected route often results in unrealistic behavior, with the adversary driving directly towards the ego vehicle in a non-human-like manner without attempting to avoid a collision.
To address these challenges, we introduce two key components: a learned objective function to select trajectories more flexibly than heuristic methods, and an adversarial skill policy that enables more human-like and reactive behavior.
Figure 1: Skill-space visualized with t-SNE [47]. Benign and adversarial priors map to several regions representing useful, human-like skills, with meaningful separation and overlap.
@@ IV-A Learned Objective Function
where $b\in\mathbb{R}$ is a hyperparameter controlling sensitivity to distance values. Both Equation 1 and Equation 2 map to $[0,1]$, where $1$ indicates maximal criticality and 00 indicates minimal. Equation 1 captures collision closeness between the ego and adversary over a given roll-out, while Equation 2 captures ego behavior difference between two episodes. However, instead of only assessing past episodes, we propose to predict these measures for a roll-out yet to happen by training a neural network, $\pi_{\text{score}}$ (detailed in Section IV-C). This $\pi_{\text{score}}$ network aims to predict $f_{\text{coll}}$ and $f_{\text{diff}}$ conditioned on a previous $\widetilde{X}^{(k)}_{\text{ego}}$ and the proposed $\widetilde{X}^{(K+1)}_{\text{adv}}$. The final score for ranking candidate trajectories is the sum of the predicted $f_{\text{coll}}$ and $f_{\text{diff}}$ values from $\pi_{\text{score}}$, averaged over the $K$ previous ego roll-outs.
@@ IV-B Adversarial Skill Learning
In our work, we separate the demonstrated skills into adversarial (i.e., those ending in a collision or near-miss) and benign skills (i.e., those avoiding a collision while staying on road). We then train two prior networks in parallel with a shared-skill VAE: benign skills flow through a “benign” prior while adversarial skills flow through an analogous “adversarial” prior. In this way, the adversarial agent policy, $\pi_{\text{adv}}$, leverages the adversarial prior to select skills which are likely to lead to safety critical outcomes, while still being a demonstrated, human-like skill, thereby improving behavior realism. Figure 1 visualizes the learned skill spaces over uniformly sampled states; regions of overlap correspond to skills which may be useful to both an adversarial and benign agent (e.g., lane-keeping, smooth kinematics, etc.) while distinct regions correspond to skills only useful for that particular agent (e.g., for an adversary: cutting-off another vehicle, hard-braking in a dangerous way, etc.).
@@ IV-C SEAL Implementation Details
For training and validating both the learned objective function and skill spaces, we leverage the well-established Waymo Open Motion Dataset (WOMD) [10] dataset, as well as a subset of scenes therein labeled by Waymo as containing interacting agents. A further subset of 500 of these scenes has been used by prior work, and we henceforth refer to this set as WOMD-Normal [1, 49]. We split these scenes into 400 training and 100 evaluation scenarios.
For $\pi_{\text{gen}}$, we utilize a pre-trained DenseTNT [13] trajectory prediction model, as used by CAT. We use the MetaDrive simulator [50] and its included IDM policy [51] as the heuristic reactive agent to collect imperfect demonstration data, described and utilized in both Section IV-A and Section IV-B. For data augmentation, all agents in the scene follow the IDM policy and produce useful demonstrations, rather than collecting examples from solely the ego.
@@ V Experimental Setup
TABLE I: Ego performance on adversarially-perturbed (a, b, c) and unmodified, real-world (d, e) scenarios. WOMD-Normal are WOMD [10] scenes with basic interactive agents labeled by Waymo; WOMD-SafeShift-Hard refers to SafeShift-mined [18] real scenes in WOMD. Adversarially-perturbed scenes use WOMD-Normal as base scenarios, in both training and evaluation settings. Higher success rates and lower crash and out of road rates are better.
@@
(d) WOMD-Normal, Real Scenes
@@
(e) WOMD-SafeShift-Hard, Real Scenes
@@
TABLE II: Scenario generation quality; results are averaged over all tested ego models. WD measures are Wasserstein distances over adversary behavior; a lower value indicates greater realism. A lower ego Success is better, as this table assesses the effectiveness of safety-critical scenario generation.
Eval Scenario Type Ego Success ($\downarrow$) Realism WD ($\downarrow$) Yaw WD ($\downarrow$) Acc WD ($\downarrow$) Road WD ($\downarrow$)
WOMD-Normal, Real Scenes 0.60 0.06 0.12 0.02 0.03
WOMD-SafeShift-Hard, Real Scenes 0.41 0.07 0.12 0.04 0.04
WOMD-Normal, GOOSE-Gen [2] 0.43 0.40 0.12 0.60 0.48
WOMD-Normal, CAT-Gen [1] 0.30 0.17 0.12 0.31 0.07
WOMD-Normal, SEAL-Gen 0.31 0.11 0.12 0.16 0.05
We leverage SEAL to generate scenes for two primary purposes: providing data augmentation during closed-loop training of reinforcement learning (RL) agent policies, and providing a means of evaluating such agents’ capabilities.
@@ V-A Policy Training
We utilize ReSkill [48] as our underlying RL algorithm, a recent SOTA approach in hierarchical RL. We use our skill-space built in Section IV-B, utilizing the benign prior rather than the adversarial one. The action learned by the ReSkill agent is a remediating $\Delta\mathbf{a}$ adjustment to the action decoded based on the current skill and state pair, $\mathbf{a^{\prime}}$. Thus, the action sent to the environment is $\mathbf{a}=\mathbf{a^{\prime}}+\Delta\mathbf{a}$. Actions are performed at a 10Hz rate, and all agents are trained for one million timesteps in total, empirically sufficient for consistent policy convergence.
@@ V-B Evaluation Settings
Many previous works evaluate agent performance, in-distribution, on a held-out subset of their own generated scenes [1, 28, 26, 34]. For additional comprehensiveness, we propose to utilize a recent scenario characterization approach, SafeShift [18], for identifying real-world safety-relevant base scenarios, denoted as WOMD-SafeShift-Hard. We start by identifying scenarios containing interacting agents labeled by Waymo. We then apply SafeShift’s hierarchical scoring to these agents and select scenarios where the interacting agents have trajectory scores in the top 20th percentile across WOMD, randomly sampling 100 scenes therein. The ego and adversary agents are assigned to the interacting agents with the higher and lower trajectory score, respectively.

### v3

@@ Abstract
Verification and validation of autonomous driving (AD) systems and components is of increasing importance, as such technology increases in real-world prevalence. Safety-critical scenario generation is a key approach to robustify AD policies through closed-loop training. However, existing approaches for scenario generation rely on simplistic objectives, resulting in overly-aggressive or non-reactive adversarial behaviors. To generate diverse adversarial yet realistic scenarios, we propose SEAL, a scenario perturbation approach that leverages learned objective functions and adversarial, human-like skills. SEAL-perturbed scenarios are more realistic than SOTA baselines, leading to improved ego task success across real-world, in-distribution, and out-of-distribution scenarios, of more than 20%. To facilitate future research, we release our code and tools: https://navars.xyz/seal/
Index Terms:
Intelligent Transportation Systems, Autonomous Vehicle Navigation, Performance Evaluation and Benchmarking
@@ I Introduction
Figure 1: An overview of SEAL. Our scenario generation approach leverages a learned objective function and an adversarial skill-based, reactive policy, for improved adversary realism and more effective closed-loop training, leading to safer autonomous driving agents, compared to previous approaches such as CAT [1] and GOOSE [2].
@@
Therefore, in this paper, we propose and evaluate a method for Skill-Enabled Adversary Learning (SEAL), which yields significantly improved downstream ego behavior, in closed-loop training with safety-critical scenario generation. Our method addresses the identified limitations in prior art by introducing two novel components, as shown in Figure 1. First, we introduce a learned objective function to anticipate how a reactive ego agent will respond to a candidate adversarial agent behavior. We quantify both collision closeness and induced ego behavior deviation, thus providing a broadened understanding of safety criticality. Second, we develop a skill-enabled, reactive adversary policy; in particular, inspired by human cognition, we leverage a hierarchical framework that is akin to how humans operate vehicles [27] and we create an adversarial prior that selects human-like skill primitives to increase criticality while maintaining realism.
Furthermore, we argue that safety-critical scenario generation should be evaluated based on behavior realism and usefulness for ego policy improvement, not just induced criticality. Prior work often assesses ego policies on generated scenarios where safety-critical behavior remains effectively in-distribution with respect to training data and heuristic perturbations [1, 28]. To address this, we build on recent scenario characterization work, SafeShift [18], to identify real (non-generated) but safety-relevant scenarios, enabling a more realistic, out-of-distribution evaluation. While in-distribution performance is informative, real-world performance on challenging scenarios is ultimately most important.
@@
    We propose two novel techniques for safety-critical perturbation: (i) a learned objective function to select candidate trajectories; and (ii) an adversarial skill-based, reactive policy for more realism in adversary behavior.
@@
    We design an improved evaluation setting for closed-loop training, utilizing real-world safety-relevant scenarios in contrast to just in-distribution generated scenarios.
@@ II-A Scenario Generation in Autonomous Driving
Approaches for generating scenarios that reproduce the distribution of normal driving behavior have been extensively explored. Some methods ensure the diversity of generated traffic behavior [29, 30], while others aim for controllability through rule-based or language-driven specifications [31, 32, 33]. However, due to the rarity of safety-critical events in recorded data [17, 18, 19], other approaches have focused on directly generating corner-case scenarios by injecting adversarial behaviors. Earlier works in safety-critical scenario generation relied on gradient-based optimization approaches with access to vehicle dynamics [28, 25, 34], a limitation in model-free settings. Other methods, such as diffusion-based approaches [26, 35], are compute-intensive and impractical to be used in a closed-loop manner. Efficient methods like CAT [1] and GOOSE [2], which leverage trajectory prediction priors and reinforcement learning (RL) respectively, prioritize simple collision objectives and are non-reactive to the ego agent. Similarly, [36] employs reactive adversaries but focuses only on collisions for criticality and defines realism via proximity to ground-truth trajectories, making it sensitive to distribution shifts. In contrast, our approach efficiently generates reactive, nuanced adversarial behavior across multiple axes of criticality, providing a stronger closed-loop training signal.
@@ II-B Robust Training and Evaluation in Autonomous Driving
Out-of-distribution evaluation has been well-explored in AD trajectory prediction [18, 43, 44, 16], but these approaches often aim to characterize an entire scenario without focusing on a single ego driver or identifying a specific adversary. In AD control tasks, some prior work has explored out-of-distribution settings, such as CARNOVEL [45, 46], which tests unseen scenario types like roundabouts. Additionally, Lu et al. [47] evaluate across real-world scenarios of various difficulty levels, but do not hold out the hardest scenarios during training. Our approach thus addresses this gap by offering a more comprehensive and rigorous evaluation, across a wide set of adversarial and real-world scenarios.
@@ III Preliminaries
Scenario Perturbation Task: For this task, $K$ re-simulations of a base scenario $\mathbf{S}$ are performed as episodes, where agents start from the same state as the base scenario and follow a behavior prescribed by some policy (i.e., a reactive policy or predefined trajectory), which may be different than their original trajectory. Let $\widetilde{\mathbf{X}}^{(k)}$ represent the observed trajectories in the $k$-th re-simulation of $\mathbf{S}$. The perturbation-based safety critical scenario generation task is thus assigning behaviors to roll-out for all non-ego agents, conditioned on the base scenario $\mathbf{S}$ and $K$ previous episodes, $\{\widetilde{\mathbf{X}}^{(k)}\ |\ k\in\{1,2,...,K\}\}$, such that the resulting $\widetilde{\mathbf{X}}^{(K+1)}$ satisfies some specified desired properties of criticality. Importantly, we treat the ego agent’s behavior as a black box: while we are able to observe previous behavior as $\widetilde{\mathbf{X}}^{K}_{\text{ego}}$, we have no access to the model or any privileged information on ego’s decision-making process. In practice, during training, we maintain a queue of the most recent $K$ perturbation roll-outs for each base scenario; during evaluation, we instead run $K$ sequential perturbation–simulation steps and use the final roll-out as the adversarial scenario.
@@ IV Approach: Skill-Enabled Adversary Learning for Scenario Generation
To increase scenario criticality while preserving realism, we propose the Skill-Enabled Adversary Learning (SEAL) approach for perturbation-based scenario generation. Similar to CAT [1], SEAL employs a probabilistic trajectory predictor $\pi_{\text{gen}}$ to sample candidate adversary futures conditioned on $\mathbf{S}$. However, directly executing these samples has three main limitations: 1) it measures criticality only via collisions, ignoring, e.g., forced ego hard braking or swerving; 2) it prevents reactivity to ego decisions; and 3) it often generates non-human-like behavior, driving straight at the ego with no avoidance. SEAL addresses these with a learned objective for flexible trajectory selection and an adversarial skill policy for more human-like and reactive behavior.
Figure 2: Skill space visualized with t-SNE [48]. Benign and adversarial priors map to several regions representing useful, human-like skills, with meaningful separation and overlap.
@@ IV-A Learned Objective Function
where $b\in\mathbb{R}$ is a hyperparameter controlling sensitivity to distance values. Both Equation 1 and Equation 2 map to $[0,1]$, where $1$ indicates maximal criticality and 00 indicates minimal. Equation 1 captures collision closeness between the ego and adversary over a given roll-out, while Equation 2 captures ego behavior difference between two episodes. However, instead of only assessing past episodes, we propose to predict these measures for a roll-out yet to happen by training a neural network, $\pi_{\text{score}}$ (detailed in Section IV-C). This $\pi_{\text{score}}$ network aims to predict $f_{\text{coll}}$ and $f_{\text{diff}}$ conditioned on a previous $\widetilde{X}^{(k)}_{\text{ego}}$ and the proposed $\widetilde{X}^{(K+1)}_{\text{adv}}$. The final score for ranking candidate trajectories is the sum of the predicted $f_{\text{coll}}$ and $f_{\text{diff}}$ values from $\pi_{\text{score}}$, averaged over the $K$ previous ego roll-outs.By using $\pi_{\text{score}}$ in place of additional heuristic simulation, we enable scoring to be conditioned on actual ego policy roll-outs and substantially reduce runtime overhead.
@@ IV-B Adversarial Skill Learning
In our work, we separate the demonstrated skills into adversarial (i.e., those ending in a collision or near-miss) and benign skills (i.e., those avoiding a collision while staying on road). We use a sliding-window partitioning scheme that excludes segments starting within twice the skill horizon before an out-of-road event, and labels as adversarial those within the same window before a collision. This reflects the intuition that not only the final skill but also preceding behavior contributes to unsafe outcomes. We then train two prior networks in parallel with a shared-skill VAE: benign skills flow through a “benign” prior while adversarial skills flow through an analogous “adversarial” prior. In this way, the adversarial agent policy, $\pi_{\text{adv}}$, leverages the adversarial prior to select skills likely to lead to safety-critical outcomes. Furthermore, because each prior is implemented as a real-valued non-volume preserving transformation trained on observed data (as in ReSkill), sampled noise vectors bijectively correspond to plausible, in-distribution behaviors. Figure 2 visualizes the learned skill spaces over uniformly sampled states; regions of overlap correspond to skills which may be useful to both an adversarial and benign agent (e.g., lane-keeping, smooth kinematics, etc.) while distinct regions correspond to skills only useful for that particular agent (e.g., for an adversary: cutting-off another vehicle, hard-braking in a dangerous way, etc.).
@@ IV-C SEAL Implementation Details
For training and validating both the learned objective function and skill spaces, we leverage the well-established Waymo Open Motion Dataset (WOMD) [10] dataset, as well as a subset of scenarios therein labeled by Waymo as containing interacting agents. A further subset of 500 of these scenarios has been used by prior work, and we henceforth refer to this set as WOMD-Normal [1, 50]. We split these scenarios into 400 training and 100 evaluation examples.
For $\pi_{\text{gen}}$, we utilize a pre-trained DenseTNT [13] trajectory prediction model, as used by CAT. $\pi_{\text{gen}}$ takes as input the first one second of X, as well as the static meta information M, and produces 32 candidate eight-second future adversary paths. We use the MetaDrive simulator [51] and its included IDM policy [52] as the heuristic reactive agent to collect imperfect demonstration data, described and utilized in both Section IV-A and Section IV-B. For data augmentation, all agents in the scenario follow the IDM policy and produce useful demonstrations, rather than collecting examples from solely the ego. We extract subgoals from each trajectory using MetaDrive’s default waypoint logic, placing checkpoints every 8 meters as navigation input for $\pi_{\text{adv}}$.
@@ V Experimental Setup
TABLE I: Ego performance on adversarially-perturbed (a, b, c) and unmodified, real-world (d, e) scenarios. WOMD-Normal are WOMD [10] scenarios with basic interactive agents labeled by Waymo; WOMD-SafeShift-Hard refers to SafeShift-mined [18] real scenarios in WOMD. Adversarially-perturbed scenarios use WOMD-Normal as base scenarios, in both training and evaluation settings. Higher success rates and lower crash and out of road rates are better. Ego realism scores are shown in (f), averaged over settings (a–e) using Wasserstein distance (WD); lower is better.
@@
(d) WOMD-Normal, Real Scenarios
@@
(e) WOMD-SafeShift-Hard, Real Scenarios
@@
(f) Aggregate Realism
Training Yaw WD Acc WD Road WD None (Replay) 0.014 0.269 0.004 No Adv 0.147 3.041 0.252 GOOSE 0.152 3.052 0.312 CAT 0.154 3.050 0.374 SEAL 0.146 3.074 0.270
TABLE II: Scenario generation quality; results are averaged over all tested ego models. WD measures are Wasserstein distances over adversary behavior; a lower value indicates greater realism. Lower collision velocities (m/s) and head-on rates are better. A lower ego Success is better, as this table assesses safety-critical effectiveness.
Eval Scenario Type Ego Success ($\downarrow$) Realism WD ($\downarrow$) Yaw WD ($\downarrow$) Acc WD ($\downarrow$) Road WD ($\downarrow$) Coll. Vel. ($\downarrow$) Head-On ($\downarrow$) Head-On (Severe) ($\downarrow$)
WOMD-Normal, Real Scenarios 60.0% 0.056 0.120 0.020 0.027 2.849 06.7% 04.2%
WOMD-SafeShift-Hard, Real Scenarios 41.3% 0.069 0.116 0.044 0.043 2.206 00.0% 00.0%
WOMD-Normal, GOOSE-Gen 43.0% 0.401 0.124 0.601 0.482 4.744 13.2% 11.7%
WOMD-Normal, CAT-Gen 29.6% 0.167 0.123 0.305 0.074 4.136 07.1% 05.9%
WOMD-Normal, SEAL-Gen 31.9% 0.108 0.121 0.157 0.049 2.950 09.2% 03.6%
We leverage SEAL to generate scenarios for two primary purposes: providing data augmentation during closed-loop training of reinforcement learning (RL) agent policies, and providing a means of evaluating such agents’ capabilities.
@@ V-A Policy Training
We utilize ReSkill [49] as our underlying RL algorithm, a recent SOTA approach in hierarchical RL. We use our skill space built in Section IV-B, utilizing the benign prior rather than the adversarial one. The low-level action learned by the ReSkill agent is a remediating $\Delta\mathbf{a}$ adjustment to the action decoded based on the current skill and state pair, $\mathbf{a^{\prime}}$, while the high-level action selects the noise vector to be passed to the prior. Thus, the action sent to the environment is $\mathbf{a}=\mathbf{a^{\prime}}+\Delta\mathbf{a}$. Actions are performed at a 10Hz rate, and all agents are trained for one million timesteps in total, empirically sufficient for consistent policy convergence.
@@ V-B Evaluation Settings
Many previous works evaluate agent performance, in-distribution, on a held-out subset of their own generated scenarios [1, 28, 26, 34]. For additional comprehensiveness, we propose to utilize a recent scenario characterization approach, SafeShift [18], for identifying real-world safety-relevant base scenarios, denoted as WOMD-SafeShift-Hard. We start by identifying scenarios containing interacting agents labeled by Waymo. We then apply SafeShift’s hierarchical scoring to these agents and select scenarios where the interacting agents have trajectory scores in the top 20th percentile across WOMD, randomly sampling 100 scenarios therein. The ego and adversary agents are assigned to the interacting agents with the higher and lower trajectory score, respectively.